How much heat energy must be added to a 6.0-cm-diameter copper sphere to raise its temperature from-50° C to 190° C? Express your answer with the appropriate units. Units
Answer #1

Solution)

Here, mass of the sphere

M= PV=8940 kg/m^3 * (4/3*3.14*(6*10^-2/2)^2)=1.0105776 kg

Now, Q= mc (T2-T1)

Q=1.0105776 * 385 * (190-(-50))= 9.3*10^4 J (Ans)

=====

Good luck!:)